OpenStudy (anonymous):

If s = 1 over 4 unit and A = 16s2, what is the value of A, in square units? ____ square unit(s). (Input whole number only.) @dtan5457 @sissyedgar

OpenStudy (sissyedgar):

I'm not sure on this one! srry

OpenStudy (dtan5457):

From what I can see. s=1/4? a=16s^2 a=16(1/4)^2

OpenStudy (holly00d1248):

1/4=.25

OpenStudy (holly00d1248):

.25*16

OpenStudy (dtan5457):

@sjman1 That's a 16s^2 right?

OpenStudy (holly00d1248):

4^2=8

OpenStudy (holly00d1248):

this help bruh

OpenStudy (anonymous):

YEAH @Holly00d1248

OpenStudy (dtan5457):

4^2 is definitely not 8 lol

OpenStudy (holly00d1248):

K bruh

OpenStudy (holly00d1248):

i mean 16 srry

OpenStudy (dtan5457):

If the question is 16s^2, it's not 16 either.

OpenStudy (dtan5457):

(order of operations)..

OpenStudy (dtan5457):

16(.25)^2=(.25^2)16=1

OpenStudy (anonymous):

so thats the awnswer @dtan5457

OpenStudy (dtan5457):

You still never confirmed if that question is 16s^2 or not. but if it is, then yeah, it is.

OpenStudy (anonymous):

yes then. @dtan5457

OpenStudy (holly00d1248):

yes the answer 1